OpenStudy (anonymous):

I need help solving for (x) on this equation!!!!!!!!!!! (x+1)/5=(2x-2)/6

OpenStudy (amigatour):

Multiply both sides by 30.

OpenStudy (amigatour):

Here is how I did it.

OpenStudy (amigatour):

30(x+1)/(5)=30(2x-2)/(6) 6(x+1) = 5(2x-2) 6x+6 = 10x - 10 -6x+10= -6x+10 16 = 4x 4 = x Check: [(4+1)]/(5) = [(2(4)-2]/(6) 5/5=6/6 1=1

OpenStudy (anonymous):

thanks